OpenStudy (anonymous):

which is true? :OOOO 1.4 < 1 1/5 1 3/20 > 1.4 2 3/5 >2.8 Orrrrrrr 2.8 < 2 17/20

jimthompson5910 (jim_thompson5910):

I would convert each number to decimal form and compare. Of course, you can convert each number to an improper fraction and compare that way, but it's up to you.

jimthompson5910 (jim_thompson5910):

Ex: 1 1/5 = 1+1/5 = 1+0.2 = 1.2 So 1.4 < 1 1/5 becomes 1.4 < 1.2 which is false